OpenStudy (mathmusician):

The Region R is bounded by the x-axis, x=1, x=3, and y=1/x^3.

OpenStudy (mathmusician):

The vertical line x = k divides the region R into two regions such that when these two regions are revolved about the x-axis, they generate solids with equal volumes. Find the value of k.

OpenStudy (mathmusician):

@FaiqRaees

OpenStudy (mathmusician):

@zepdrix

OpenStudy (mathmusician):

This is as far as I got \[\pi \int\limits_{1}^{k}(\frac{ 1 }{ x ^{2} })^{2} dx = \pi \int\limits_{k}^{3}(\frac{ 1 }{ x ^{3} })^{2}dx\]

OpenStudy (mathmusician):

both of the denominators are supposed to be x^3

zepdrix (zepdrix):

Hmm looks good so far :) Divide the pi's,\[\large\rm \int\limits\limits_{1}^{k}\frac{1}{x^6}dx = \int\limits\limits_{k}^{3}\frac{1}{x^6}dx\]then expand out the integrals and solve for k.

OpenStudy (mathmusician):

okay

OpenStudy (mathmusician):

i got here

zepdrix (zepdrix):

ok let's divide the -1/5 out of each side. It just makes it easier to plug everything in if they're not floating around.